# Learning Mean Field Control on Sparse Graphs

### v1

@@ Appendix C Simulation Details
We use MARLlib $1.0.0$ (Hu et al., 2023a) building on RLlib $1.8.0$ (Apache-2.0 license) (Liang et al., 2018) and its PPO implementation (Schulman et al., 2017) for IPPO and our algorithms. For our experiments, we used around 80 0008000080\,00080 000 core hours on CPUs, and each training run usually took a single day of training on up to $96$ parallel CPU cores. For the policies we used two hidden layers of $256$ nodes with $\tanh$ activations. We used a discount factor of $\gamma=0.99$ with GAE $\lambda=1.0$, and training and minibatch sizes of $4000$ and $1000$, performing $5$ updates per training batch. The KL coefficient and clip parameter were set to $0.2$, with a KL target of $0.03$. The learning rate was set to $0.00005$. The problem details are found in the following.

### v2

@@ Appendix C Simulation Details
We use MARLlib $1.0.0$ (hu2023marllib) building on RLlib $1.8.0$ (Apache-2.0 license) (liang2018rllib) and its PPO implementation (schulman2017proximal) for IPPO and our algorithms. For our experiments, we used around 80 00080\,000 core hours on CPUs, and each training run usually took a single day of training on up to $96$ parallel CPU cores. For the policies we used two hidden layers of $256$ nodes with $\tanh$ activations. We used a discount factor of $\gamma=0.99$ with GAE $\lambda=1.0$, and training and minibatch sizes of $4000$ and $1000$, performing $5$ updates per training batch. The KL coefficient and clip parameter were set to $0.2$, with a KL target of $0.03$. The learning rate was set to $0.00005$. The problem details are found in the following.